OpenStudy (anonymous):

The sum of two numbers is 12 find the minimum value of the sum of their cubes a. 432 b. 346 c. 644 d. 244

OpenStudy (anonymous):

5 and 7 is near lol =244

OpenStudy (anonymous):

3 and 9 is near to =243 <smaller 1

OpenStudy (anonymous):

x+y = 12 p(x)= x^3 + y^3 = (x+y)^3 - 3xy(x+y) p(x) =12^3 -36xy p(x)=1728 -36xy minimise p(x) like we did last time set p'(x) = 0

OpenStudy (anonymous):

first substitute y = 12 -x

OpenStudy (anonymous):

the key to making this problem easier was the algebraic identity (x+y)^3 -3xy(x+y) = x^3 + y^3

OpenStudy (anonymous):

yup i heard early its positive but if the answer is true to be negative it must.

OpenStudy (anonymous):

oooh er i think i made a mistake somewhere..

OpenStudy (anonymous):

p(x)=1728 -36xy p(x)=1728 -36x(12 - x) = 1728 - 432x + 36x^2 p'(x) = 72x - 432 set p'(x) = 0 x = 432/72 = 6 y = 12-6 = 6

OpenStudy (anonymous):

so 6^3 + 6^3 = 432

OpenStudy (anonymous):

the answer is either 6,6 or 1,11 isn't it? boundary problem, special values... (6,6) = 432 (1,11) = 1,331 so I will go with 432

OpenStudy (anonymous):

and I was wrong, the answer was going to be either 6,6 or 0,12

OpenStudy (anonymous):

i solved it using a bit of calculus

OpenStudy (anonymous):

eigen if this question is question again but the numbers are different how could i get like 12^3? in fastest way

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

i get it

OpenStudy (anonymous):

\[(x+y)^3 = x^3 + 3x^2y + 3xy^2 + y^3\] rearrange to get \[x^3 + y^3\] \[x^3 + y^3 = (x+y)^3 - 3x^2y - 3xy^2 = (x+y)^3 - 3xy(x + y)\] this is useful because we know what x+y is equal to :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem, happy to help